# Materialist: Physically Based Editing Using Single-Image Inverse Rendering

### v1

@@ 8.1 Optimize SH coefficients instead of envmap
Table 5: Time taken by different methods to optimize $\mathcal{L}_{\text{re}}$ to 0.03. Direct envmap optimization is faster than optimize SH coefficient.
@@
8.2 Optimization Steps and Rerendering error
We analyzed the relationship between optimization steps and rerender error. We randomly selected 10 images from the InteriorVerse dataset [78] to test rerendering error at different optimization steps. Optimization on an RTX 3090 runs at 2 steps per second. Based on MatNet’s accuracy, satisfactory results are typically achieved after 10 minutes of optimization. Results are shown in Table. 6.
Table 6: Ablation of Optimization steps. Compare the rerender error.
@@ 8.3 Optimization Network
Figure 13: Compared to a CNN-based UNet, the position-embedded MLP achieves faster optimization.
@@
Figure 14: UNet Artifect. The red box highlights an enlarged view of this area, and the details is in the second row.
@@
Table 7: Comparison of the lowest $\mathcal{L}_{\text{re}}$ that can be obtained using MLP and directly optimizing material properties when the predicted roughness and metallic are not provided.
@@ 8.4 Optimizing Normal
Figure 15: Optimizing normal at the same time will make the optimization process more unstable.
9 Limitations
In this section, we discuss potential limitations of our proposed method from a broader, user-oriented perspective. It is important to note that some points, such as the need for per-image optimization, are not typically viewed as limitations within the differentiable rendering field. However, from a user’s viewpoint, this may make our approach less convenient than stable diffusion-based methods [53, 7].
1. As with all differentiable rendering methods, our approach requires optimization for each image to obtain the lighting and optimized material properties, which might be quite time-consuming. Depending on MatNet’s prediction accuracy, optimization takes between 5 to 30 minutes. For out-of-domain images (i.e., non-indoor scenes) or higher-quality requirements, optimization time may increase. This limits our method’s ability to achieve the rapid, batch image editing enabled by stable diffusion-based techniques [7]. Training MatNet on a broader dataset for more accurate predictions could help mitigate this limitation.
4. For Material Transparency Editing, complex object geometry can lead to inaccurate refraction distortions, as our method assumes only two refractions. Single-view mesh reconstruction [31, 32, 54] may yield more accurate refractions for complex shapes. Besides, since we simulate refraction without the rays truly passing through the object, shadows cast by transparent objects may lack accuracy, and does not have caustic effect. (Note that this problem does not exist for object insertion tasks with complete geometry.)
Figure 16: Super-resolution often struggles with refraction effects. As shown, both Super Resolution 1 and Super Resolution 2 exhibit noticeable refraction issues, requiring multiple runs to achieve a satisfactory result, as seen in Super Resolution 3.
10 Refraction Length Prediction
@@ 10.1 Dataset Preparation
$d=\begin{cases}\|\mathcal{P}_{n+1}-\mathcal{P}_{n}\|,&\text{if }\vec{\omega}_{r}\cdot\vec{n}>0\text{ \& hollow object}\\ \|\mathcal{P}_{n+1}-\mathcal{P}_{n}\|,&\text{if }\vec{\omega}_{r}\cdot\vec{n}<0\text{ \& non-hollow object}\\ 0,&\text{Otherwise }\end{cases}$ (24)
@@
Furthermore, "hollow" objects represent things like water jugs with air in between the surfaces and "non-hollow" objects are things like lemon with no air inside the object.
@@
Figure 17: Example of refraction length dataset. The closer to white, the longer the refraction distance.
11 Material Editing Settings
@@
12 Material Editing User Study
We conducted a user study on material editing results, asking participants to rank the editing results from Alchemist [53], ours (direct rendering), and ours SR (rendered result refined with Super Resolution). The highest rank received 3 points, and the lowest 1 point. Since these are real-world photos with no ground truth, participants were provided with synthetic scene renderings as reference and asked to imagine the expected edit on real photos as closely as possible. It should be noted that participants generally found it challenging to envision the correct editing result, making it difficult to judge which method was better. When uncertain, participants often zoomed in to examine details for ranking, which may have led to lower scores for our direct rendering. However, our scores improved significantly after applying super resolution.
Table 8: Material Editing User Study. SR represent for super resolution.
Edit Task Alchemist [53] Ours Ours (SR)
@@ 12 Material Editing User Study
Figure 18: Our methods with super-resolution. Using super resolution (SR) can greatly improve the quality of the rendered image. However, the current SR model cannot fully retain the details of the original image and will make excessive modifications to the original image. We also provide the effect of modifications to material properties on appearance in a synthetic dataset.
13 Coordinate Transformation
@@
$\mathbf{N}=\bm{P}\cdot\mathbf{Q}$ (26)
@@
$x_{\text{screen}}=\left(\frac{x_{\text{ndc}}+1}{2}\right)\cdot W$ (27)

### v2

@@ B.1. Optimize SH coefficients instead of envmap
Table 5. Time taken by different methods to optimize $\mathcal{L}_{\text{re}}$ to 0.03. Direct envmap optimization is faster than optimize SH coefficient.
@@
B.2. Optimization Steps and Rerendering error
We analyzed the relationship between optimization steps and rerender error. We randomly selected 10 images from the InteriorVerse dataset (Zhu et al., 2022b) to test rerendering error at different optimization steps. Optimization on an RTX 3090 runs at 2 steps per second. Based on MatNet’s accuracy, satisfactory results are typically achieved after 10 minutes of optimization. Results are shown in Table. 6.
Table 6. Ablation of Optimization steps. Compare the rerender error.
@@ B.3. Optimization Network
Figure 17. Compared to a CNN-based UNet, the position-embedded MLP achieves faster optimization.
@@
Figure 18. UNet Artifect. The red box highlights an enlarged view of this area, and the details is in the second row.
@@
Table 7. Comparison of the lowest $\mathcal{L}_{\text{re}}$ that can be obtained using MLP and directly optimizing material properties when the predicted roughness and metallic are not provided.
@@ B.4. Optimizing Normal
Figure 19. Optimizing normal at the same time will make the optimization process more unstable.
Appendix C Simplified DisneyBRDF
To model surface scattering we use the DisneyBRDF (Burley, 2012, 2015) for our implementation. During the optimization process, we do not model sheen, clearcoat, and glass, so the BRDF in Eq. 4 can be written
(17) $f_{r}(\bm{x},\vec{\omega}_{i},\vec{\omega}_{o})=f_{s}+(1-\mathbf{M}(\bm{x}_{s}))f_{d}$
with the specular ($f_{s}$) and diffuse ($f_{d}$) terms defined by
(18) $f_{s}$ $=\frac{F_{s}(\vec{\omega}_{i},\vec{\omega}_{o};\eta)D_{s}(\vec{h};\mathbf{R}(\bm{x}_{s}))G_{s}(\vec{\omega}_{i},\vec{\omega_{o}};\mathbf{R}(\bm{x}_{s}))}{4\,|\vec{n}\cdot\vec{\omega}_{i}||\vec{n}\cdot\vec{\omega}_{o}|}$
(19) $f_{d}$ $=\frac{\mathbf{A}(\bm{x}_{s})}{\pi}F_{d}(\vec{\omega}_{i})F_{d}(\vec{\omega}_{o})\,,$
where $F$ is the Fresnel reflectance, which here uses the half vector $\vec{h}=(\vec{\omega}_{i}+\vec{\omega}_{o})/\|\vec{\omega}_{i}+\vec{\omega}_{o}\|$ and models the fraction of light reflected from the surface as a function of material properties, including the relative index of refraction $\eta$. We use the modified Schlick approximation (Burley, 2012):
(20) $\!\!F_{s}(\vec{\omega}_{i},\vec{\omega}_{o};\eta)$ $=C_{0}(\eta)+(1-C_{0}(\eta))(1-\vec{h}\cdot\vec{\omega}_{o})^{5}\,$
(21) $C_{0}(\eta)$ $=R_{0}(\eta)(1-\mathbf{M}(\bm{x}_{s}))+\mathbf{M}(\bm{x}_{s})\mathbf{A}(\bm{x}_{s})\,,$
(22) $F_{d}(\vec{\omega})$ $=1+(F_{D90}-1)(1-\vec{n}\cdot\vec{\omega})^{5}\,$
(23) $F_{D90}$ $=0.5+2\,\mathbf{R}(\bm{x}_{s})\,(\vec{h}\cdot\vec{\omega})^{2}\,.$
Regarding the other terms, $D_{s}$ is the microfacet normal distribution function, while $G_{s}$ is the microfacet shadowing-masking function (Schlick, 1994). We use the GGX distribution (Walter et al., 2007) for these terms due to its simplicity, with just one roughness parameter $\mathbf{R}(\bm{x}_{s})$, while being able to fit empirical data well.
Appendix D Refraction Length Prediction
@@ D.1. Dataset Preparation
(24) $d=\begin{cases}\|\mathcal{P}_{n+1}-\mathcal{P}_{n}\|,&\text{if }\vec{\omega}_{r}\cdot\vec{n}>0\text{ \& hollow object}\\ \|\mathcal{P}_{n+1}-\mathcal{P}_{n}\|,&\text{if }\vec{\omega}_{r}\cdot\vec{n}<0\text{ \& non-hollow object}\\ 0,&\text{Otherwise }\end{cases}$
@@
Furthermore, ”hollow” objects represent things like water jugs with air in between the surfaces and ”non-hollow” objects are things like lemon with no air inside the object.
@@
Figure 20. Example of refraction length dataset. The closer to white, the longer the refraction distance.
Appendix E Material Editing Settings
@@
Appendix F Material Editing User Study
We conducted a user study on material editing results with 34 participants, who were recruited via email from the authors’ institution. We asked participants to rank the editing results from Alchemist (Sharma et al., 2024), ours (direct rendering), and ours SR (rendered result refined with Super Resolution). The highest rank received 3 points, and the lowest 1 point. Since these are real-world photos with no ground truth, participants were provided with synthetic scene renderings as reference and asked to imagine the expected edit on real photos as closely as possible. It should be noted that participants generally found it challenging to envision the correct editing result, making it difficult to judge which method was better. When uncertain, participants often zoomed in to examine details for ranking, which may have led to lower scores for our direct rendering. However, our scores improved significantly after applying super resolution.
Table 8. Material Editing User Study. SR represent for super resolution.
Edit Task Alchemist (Sharma et al., 2024) Ours Ours (SR)
@@ Appendix F Material Editing User Study
Figure 21. Our methods with super-resolution. Using super resolution (SR) can greatly improve the quality of the rendered image. However, the current SR model cannot fully retain the details of the original image and will make excessive modifications to the original image. We also provide the effect of modifications to material properties on appearance in a synthetic dataset.
Appendix G Coordinate Transformation
@@
(26) $\mathbf{N}=\bm{P}\cdot\mathbf{Q}$
@@
(27) $x_{\text{screen}}=\left(\frac{x_{\text{ndc}}+1}{2}\right)\cdot W$
@@
Appendix H Handling High Dynamic Range (HDR) Images
High Dynamic Range (HDR) images, characterized by their wide range of luminance levels, pose stability challenges during the differentiable rendering optimizations. The extensive variation in pixel intensities can lead to unstable gradients, making optimization difficult. To mitigate this issue, a common practice is to clip the image intensity values to a predefined range, typically $[0,10]$. This clipping helps to limit the extreme values that can cause numerical instability during backpropagation. In cases where the images are extremely bright, additional preprocessing steps are often necessary. One effective approach is to apply a logarithmic transform, specifically $\log(1+x)$, to the image data. This transformation compresses the dynamic range of the image, reducing the disparity between the brightest and darkest regions. By doing so, the gradients become smoother and more manageable, facilitating more stable and efficient optimization. The $\log(1+x)$ transform is particularly useful because it preserves the relative differences in intensity while compressing the overall range, ensuring that the model can learn effectively from the data without being overwhelmed by extreme values.